An inventor claims to have developed an engine which takes in 11times10^(7)J at 400K rejects 5times10^(7)J at 200K and delivers 16.67kW hours of work.Would you advise investing money in this project?
1
Expert's answer
2021-04-13T06:36:23-0400

Solution.

"Q_1=11\sdot10^7J;"

"T_1=400K;"

"Q_2=5\sdot10^7J;"

"T_2=200K;"

"\eta=1-\dfrac{T_1}{T_2};"

"\eta=1-\dfrac{400K}{200K}=0.5 or 50\%" ;

"\eta=\dfrac{Q_1-Q_2}{Q_1};"

"\eta=\dfrac{11\sdot10^7J-5\sdot10^7J}{11\sdot10^7J}=0.55 or 55\%" ;

By Carnot's theorem, no engine can be more efficient than a carnot engine, or: Efficiency(carnot) > Efficiency(hypothetical)​and in this case, it doesnt hold, i.e.: 50 % > 55 % is not true.​

Answer: You should not invest in such a project, it is not real.